The Life of an electric component has an exponential distribution with a mean of 5 years. What is the probability that a randomly selected one such component has a life less than 5 years?  (keep 4 decimal places)
Answer #1

X ~ Exp(\lambda)

Where , Mean = 1/ \lambda = 5

So, \lambda = 1 / 5

For exponential distribution,

P(X < x) = 1 - e-\lambdaX

So,

P( X < 5) = 1 - e-5/5

= 1 - e-1

= 0.6321